### v1

# Faster Online Computation of String Net Frequency

Abstract
An occurrence of a repeated substring $u$ in a string $S$ is called a net occurrence if extending the occurrence to the left or to the right decreases the number of occurrences to 1. The net frequency (NF) of a repeated substring $u$ in a string $S$ is the number of net occurrences of $u$ in $S$. Very recently, Guo et al. [SPIRE 2024] proposed an online $O(n\log\sigma)$-time algorithm that maintains a data structure of $O(n)$ space which answers Single-NF queries in $O(m\log\sigma+\sigma^{2})$ time and reports all answers of the All-NF problem in $O(n\sigma^{2})$ time. Here, $n$ is the length of the input string $S$, $m$ is the query pattern length, and $\sigma$ is the alphabet size. The $\sigma^{2}$ term is a major drawback of their method since computing string net frequencies is originally motivated for Chinese language processing where $\sigma$ can be thousands large. This paper presents an improved online $O(n\log\sigma)$-time algorithm, which answers Single-NF queries in $O(m\log\sigma)$ time and reports all answers to the All-NF problem in output-optimal $O(|\mathsf{NF^{+}}(S)|)$ time, where $\mathsf{NF^{+}}(S)$ is the set of substrings of $S$ paired with their positive NF values. We note that $|\mathsf{NF^{+}}(S)|=O(n)$ always holds. In contract to Guo et al.’s algorithm that is based on Ukkonen’s suffix tree construction, our algorithm is based on Weiner’s suffix tree construction.
1 Introduction
[…]
Recently, Guo, Umboh, Wirth, and Zobel [9] considered the online version of the problem of computing net frequency, where the characters of the input string arrive one by one. The GUWZ algorithm given in [9] uses the suffix tree [17] of $O(n)$ space and answers Single-NF queries in $O(m\log\sigma+\sigma^{2})$ time [^2] The alphabet size $\sigma$ is assumed to be a constant in the literature [8, 9].. For the online All-NF problem, the GUWZ algorithm maintains a data structure for an input online string $S$ that takes $O(n\sigma^{2})$ time to report the answers to the All-NF problem. The GUWZ algorithm is based on Ukkonen’s left-to-right online suffix tree construction algorithm [16] and uses $O(\log\sigma)$ amortized time for updates per new character.
There are two major drawbacks in the GUWZ algorithm. The first drawback is its dependency in the alphabet size. The $\sigma^{2}$ terms in the $O(m\log\sigma+\sigma^{2})$ query time for the Single-NF problem and in the $O(n\sigma^{2})$ reporting time for the All-NF problem can be a major bottleneck in Chinese language text processing in which the alphabet size $\sigma$ is thousands large. The second drawback is that the GUWZ algorithm heavily relies on the suffix-extension data structure of Breslauer and Italiano [4], which further relies on a nearest marked ancestor (NMA) data structure [18] on dynamic trees and an order-maintenance data structure [15, 6, 1] on dynamic lists.
In this paper, we present an improved algorithm for the Single-NF and All-NF problems in the online setting (see also Table 1). We present an online algorithm that uses $O(n)$ space and maintains a list of substrings with positive NF values in $O(\log\sigma)$ amortized time per character, for general ordered alphabets of size $\sigma$. Since the positive NF values can be stored in the corresponding nodes in the suffix tree with $O(n)$ space, our data structure answers Single-NF queries in $O(m\log\sigma)$ time each. Also, our algorithm maintains a data structure that can report all answers to the All-NF problem in output optimal time, independently of the alphabet size $\sigma$.
Our algorithm is based on Weiner’s suffix tree construction [17], that works in a right-to-left online manner. For ease of explanations, we use the version of Weiner’s algorithm that uses soft Weiner links on the suffix tree, which are secondary edges of the directed acyclic word graph (DAWG) [2] of the reversed string.
Our online NF computation algorithm is built on deep insights to how Weiner’s online suffix tree construction algorithm maintains structural changes in the suffix tree in the online setting. Notably, no special data structures or operations are needed for computing the string NFs online - all the necessary information is computed by the classical Weiner algorithm and the string NFs can be obtained almost for free during the online computation of the suffix tree. We note that our running time $O(n\log\sigma)$ is optimal as long as we maintain a data structure based on suffix trees in the online string [7].
Table 1: Comparison of algorithms for computing string net frequencies for input string $S$ of length $n$ over an alphabet of size $\sigma$. $\mathsf{NF^{+}}(S)$ denotes the set of substrings of $S$ paired with their positive NF values, where $|\mathsf{NF^{+}}(S)|=O(n)$ always holds. The $\log\sigma$ factors in the construction times of the offline algorithms [8, 14] can be removed for integer alphabets of size $\sigma=n^{O(1)}$.
algorithm construction time space Single-NF time All-NF time model
GEWZ [8] $O(n\log\sigma)$ $O(n)$ $O(m+\log n+\sigma)$ $O(n)$ offline
OBO [14] $O(n\log\sigma)$ $O(n)$ $O(n)$ offline
GUWZ [9] $O(n\log\sigma)$ $O(n)$ $O(m\log\sigma+\sigma^{2})$ $O(n\sigma^{2})$ online
Our Algorithm $O(n\log\sigma)$ $O(n)$ $O(m\log\sigma)$ $O(|\mathsf{NF^{+}}(S)|)$ online
2 Preliminaries
2.1 Strings
Let $\Sigma$ be a general ordered alphabet. Any element of $\Sigma^{*}$ is called a string. For any string $S$, let $|S|$ denote its length. Let $\varepsilon$ be the empty string, namely, $|\varepsilon|=0$. Let $\Sigma^{+}=\Sigma\setminus\{\varepsilon\}$. If $S=XYZ$, then $X$, $Y$, and $Z$ are called a prefix, a substring, and a suffix of $S$, respectively. For any $1\leq i\leq j\leq|S|$, let $S[i..j]$ denote the substring of $S$ that begins at position $i$ and ends at position $j$ in $S$. For convenience, let $S[i..j]=\varepsilon$ if $i>j$. For any $1\leq i\leq|S|$, let $S[i]$ denote the $i$th character of $S$. For any string $T$, let $\mathsf{Substr}(S)$ and $\mathsf{Suffix}(S)$ denote the sets of substrings and suffixes of $S$, respectively. For any string $S$, let $S^{R}$ denote the reversed string of $S$, i.e., $S^{R}=S[|S|]\cdots S[1]$.
2.2 String Net Frequency and Maximal Repeats
For a string $w$ and a string $S$, let $\mathsf{occ}_{S}(w)$ denote the number of occurrences of $w$ in $S$, i.e.
$\mathsf{occ}_{S}(w)=|\{i\mid S[i..i+|w|-1]=w\}|.$
A string $u$ with $\mathsf{occ}_{S}(u)\geq 2$ is called a repeat in $S$. For a repeat $u\in\mathsf{Substr}(S)$ of a string $S$, let
$\Phi_{S}(u)=\{(a,b)\in\Sigma\times\Sigma\mid\mathsf{occ}_{S}(aub)=\mathsf{occ}_{S}(au)=\mathsf{occ}_{S}(ub)=1\}.$
The net frequency (NF) of $u$ in $S$, denoted $\phi_{S}(u),$ is the number of ordered pairs $(a,b)\in\Sigma\times\Sigma$ of characters such that $\mathsf{occ}_{S}(aub)=\mathsf{occ}_{S}(au)=\mathsf{occ}_{S}(ub)=1$, i.e., $\phi_{S}(u)=|\Phi_{S}(u)|$.
A repeat $u$ in a string $S$ is called a left-maximal repeat in $S$ if $u$ is a prefix of $S$, or there are at least two distinct characters $a,a^{\prime}$ that immediately precede $u$ in $S$ (i.e. $au,a^{\prime}b\in\mathsf{Substr}(S)$). Similarly, a repeat $u$ in $S$ is called a right-maximal repeat in $S$ if $u$ is a suffix of $S$, or there are at least two distinct characters $b,b^{\prime}$ that immediately follow $u$ in $S$ (i.e. $ub,ub^{\prime}\in\mathsf{Substr}(S)$). Let $\mathsf{LM}(S)$ and $\mathsf{RM}(S)$ denote the sets of left-maximal and right-maximal repeats in $S$, respectively.
For a left-maximal repeat $u$ in $S$, a character $a\in\Sigma$ such that $\mathsf{occ}_{S}(au)\geq 1$ is called a left-extension of $u$ in $S$. Similarly, for a right-maximal repeat $u$ in $S$, a character $b\in\Sigma$ such that $\mathsf{occ}_{S}(ub)\geq 1$ is called a right-extension of $u$ in $S$.
A repeat $u$ is called a maximal repeat in $S$ if $u\in\mathsf{LM}(S)\cap\mathsf{RM}(S)$. Let $\mathsf{M}(S)=\mathsf{LM}(S)\cap\mathsf{RM}(S)$ denote the set of maximal repeats in $S$.
Guo et al. [8] showed that if $\phi_{S}(u)\geq 1$, then $u\in\mathsf{RM}(S)$. Below we present a stronger statement.
Lemma 1.
If $\phi_{S}(u)\geq 1$, then $u\in\mathsf{M}(S)$.
Proof.
Since $\phi_{S}(u)\geq 1$, there exists a pair $(a,b)$ of characters such that $\mathsf{occ}_{S}(au)=\mathsf{occ}_{S}(ub)=1$. Since $\mathsf{occ}_{S}(u)\geq 2$, there is an occurrence $u=S[i..i+|u|-1]$ of $u$ such that $S[i-1]\neq a$ or $i=1$. Similarly, there is an occurrence $u=S[j-|u|+1..j]$ of $u$ such that $S[j+1]\neq b$ or $j=|S|$. Thus $u\in\mathsf{LM}(S)\cap\mathsf{RM}(S)=\mathsf{M}(S)$. ∎
Let
$\mathsf{NF^{+}}(S)=\{(u,\phi_{S}(u))\mid u\in\mathsf{Substr}(S),\phi_{S}(u)\geq 1\}$
denote the set of pairs of string $u$ and its NF that is positive. It follows from Lemma 1 that if $(u,\phi_{S}(u))\in\mathsf{NF^{+}}(S)$, then $u\in\mathsf{M}(S)$. Since $|\mathsf{M}(S)|\leq n-1$ for any string of length $n$ (c.f. [3]), it is immediate that $|\mathsf{NF^{+}}(S)|\leq n-1$.
2.3 Online Computation of String Net Frequency
We will design an online algorithm for computing substrings that have positive values of net frequencies, in which new characters are prepended to the input string. Namely, our algorithm works in a right-to-left online manner. Note that $\phi_{S}(u)$ is closed under reversal, i.e., $\phi_{S}(u)=\phi_{S^{R}}(u^{R})$. Hence, by reversing the input string, our algorithm can be used in the left-to-right online setting as well.
We consider the following problems of computing string net frequencies online.
Problem 1 (Right-to-Left Online NF Computation).
Let $S$ be a string of length $n$ for which $\mathsf{NF^{+}}(S)$ has been computed. Given a character $a\in\Sigma$, compute $\mathsf{NF^{+}}(aS)\Delta\mathsf{NF^{+}}(S)=\left(\mathsf{NF^{+}}(aS)\setminus\mathsf{NF^{+}}(S)\right)\cup\left(\mathsf{NF^{+}}(S)\setminus\mathsf{NF^{+}}(aS)\right)$.
Problem 1 requires one to perform the following per new character:
  1. (i)
    Update the NFs of some existing substrings.
  2. (i)
    Add substrings whose NFs become positive;
  3. (iii)
    Remove substrings whose NFs become 0;
2.4 Suffix Trees and Weiner Links
For convenience, we assume that each string $S$ terminates with a unique character \$currency-dollar\\$\$ that does not appear elsewhere in $S$.
A compacted trie is a rooted tree such that
  • each edge is labeled by a non-empty string,
  • each internal node is branching, and
  • the string labels of the out-going edges of each node begin with mutually distinct characters.
The suffix tree [17] for a string $S$, denoted $\mathsf{STree}(S)$, is a compacted trie which represents $\mathsf{Suffix}(S)$. We sometimes identify node $v$ with the substring it represents.
For any string $S$ of length $n$ that terminates with \$currency-dollar\\$\$, $\mathsf{STree}(S)$ has exactly $n$ leaves, and has at most $n-1$ internal nodes (including the root) since every non-leaf node is branching. By representing each edge label $x$ with an ordered pair $\langle i,j\rangle$ of integers such that $x=S[i..j]$, $\mathsf{STree}(S)$ can be stored in $O(n)$ space.
We define the suffix link of each non-root node $av$ of $\mathsf{STree}(S)$ with $a\in\Sigma$ and $v\in\Sigma^{*}$, by $\mathsf{slink}(av)=v$. For each node $v$ and $a\in\Sigma$, we also define the reversed suffix link (a.k.a. Weiner link) by ${\mathsf{W\_link}}_{a}({v})=avy$, where $y\in\Sigma^{*}$ is the shortest string such that $avy$ is a node of $\mathsf{STree}(S)$. ${\mathsf{W\_link}}_{a}({v})$ is undefined if $av\notin\mathsf{Substr}(S)$. A Weiner link ${\mathsf{W\_link}}_{a}({v})=avy$ is said to be hard if $y=\varepsilon$, and soft if $y\in\Sigma^{+}$.
It is clear that every internal node in $\mathsf{STree}(S)$ represents a right-maximal repeat. Hence, the first characters in the out-edges of each node $u$ are right-extensions of the corresponding right-maximal repeat $u$ in $S$. Also, for each internal node $u$ that is also a left-maximal repeat (i.e., $u\in\mathsf{M}(S)$), the characters $a\in\Sigma$, for which ${\mathsf{W\_link}}_{a}({u})$ are defined, are the left-extensions of $u$ in $S$.
Lemma 2 ([2]).
For any string $S$ of length $n>2$, the total number of hard and soft Weiner links in $\mathsf{STree}(S)$ is at most $43n-$ 4.
3 Algorithm
Let $S$ be the string. As in the previous work [9], a node $u$ in $\mathsf{STree}(S)$ stores the NF value $\phi_{S}(u)$ iff $\phi_{S}(u)\geq 1$.
3.1 Key Lemma
Our algorithm is based on the following lemma:
Lemma 3 (Unique Character Pairs).
Let $u$ be a repeat in a string $S$. Let $(a,b)\in\Phi_{S}(u)$. Then, for any character $c\neq b$, $(a,c)\notin\Phi_{S}(u)$. Similarly, for any character $d\neq a$, $(d,b)\notin\Phi_{S}(u)$.
Proof.
For a contradiction assume that $(a,c)\in\Phi_{S}(u)$. Then, $\mathsf{occ}_{S}(aub)=1$ and $\mathsf{occ}_{S}(auc)=1$. Since $b\neq c$, $\mathsf{occ}_{S}(au)\geq 2$. However, since $(a,b)\in\Phi_{S}(u)$, we have $\mathsf{occ}_{S}(au)=1$, a contradiction. The case for $(d,b)$ is symmetric. ∎
Lemma 3 implies that for any character $a\in\Sigma$ that is a left-extension of a (maximal) repeat $u$ in $S$, there is at most one character $b\in\Sigma$ that contributes to the positive NF of $u$ in $S$. Thus the following corollary is immediate:
Corollary 1.
For any repeat $u$ in a string $S$, $\phi_{S}(u)\leq\min\{|\Sigma^{L}_{u}|,|\Sigma^{R}_{u}|\}$, where $\Sigma^{L}_{u}$ and $\Sigma^{R}_{u}$ are the sets of left-extensions and right-extensions of $u$ in $S$, respectively.
Recall that the right-extensions of (right) maximal repeats are the first characters of the edge labels of $\mathsf{STree}(S)$. Since any string of length $n$ can contain at most $n-2$ non-empty maximal repeats [3], the next lemma follows from Lemma 1 and Corollary 1.
Lemma 4.
For any string $S$ of length $n$, $\sum_{u\in\mathsf{Substr}(S)}\phi_{S}(u)\leq n-2$.
This is a slight improvement over the upper bound $\sum_{u\in\mathsf{Substr}(S)}\phi_{S}(u)\leq n$ given by Guo et al. [8].
3.2 Recalling Weiner’s Suffix Tree Construction
In this subsection, we briefly recall how Weiner’s algorithm updates $\mathsf{STree}(S)$ for a string $S$ to $\mathsf{STree}(aS)$ with a new character $a$.
Suppose that all the Weiner links have also been computed for the nodes of $\mathsf{STree}(S)$. Since our input string $S$ terminates with \$currency-dollar\\$\$ at its right end, all the suffixes of $S$ are represented by the leaves of $\mathsf{STree}(S)$.
To update $\mathsf{STree}(S)$ into $\mathsf{STree}(aS)$, it is required to insert a new leaf that represents the new suffix $aS$. To insert a new leaf for $aS$, Weiner’s algorithm finds the locus of the longest repeating prefix $av$ of $aS$ on $\mathsf{STree}(S)$. Then, $av$ will be the parent of the new leaf for $aS$.
Refer to caption Refer to caption
Figure 1: A snapshot of Weiner’s suffix tree construction algorithm. Solid arcs represent edges, and broken thick (resp. thin) arcs represent hard (resp. soft) Weiner links. Left: By climbing up the path from leaf $S$, we find the lowest ancestor $v$ that has a soft Weiner link ${\mathsf{W\_link}}_{a}({v})=ax$ that is labeled by $a$, and the lowest ancestor $z$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})=az$ that is labeled by $a$. Right: Split the edge from node $az$ to $ax$ at the locus for $av$, and insert a new leaf for $aS$ from $av$. Then, the NF of node $u$ is increased by 1, since $(a,b)\in\Phi_{aS}(u)\setminus\Phi_{S}(u)$.
In Weiner’s algorithm, we climb up the path from the leaf for $S$ and finds the first ancestor $v$ such that the Weiner link ${\mathsf{W\_link}}_{a}({v})$ with the new character $a$ is defined. See also Figure 1 for illustration. There are two cases:
(1) If ${\mathsf{W\_link}}_{a}({v})$ is a soft Weiner link, then let $ax$ be the destination node of this Weiner link. Note that $av$ is a proper prefix of $ax$. We continue climbing up the path until finding the lowest ancestor $z$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ with $a$. We move to node $az$ using the hard Weiner link ${\mathsf{W\_link}}_{a}({z})=az$. The locus for $av$ is on the edge from $az$ to $ax$. We split this edge at the locus for $av$, and insert the new leaf for $aS$. Since $av$ is a node in the updated suffix tree $\mathsf{STree}(aS)$, the Weiner link ${\mathsf{W\_link}}_{a}({v})$ becomes hard and points to $av$. The soft Weiner links of the nodes between $v$ and $z$ exclusive are redirected to $av$. Finally, we copy all (soft or hard) Weiner links of node $ax$ to the new node $av$, namely, if ${\mathsf{W\_link}}_{d}({ax})=w$ for character $d\in\Sigma$ and node $w=dax$, then we set ${\mathsf{W\_link}}_{d}({av})=w$.
(2) If ${\mathsf{W\_link}}_{a}({v})$ is a hard Weiner link, then we move to node $av$ using the hard Weiner link ${\mathsf{W\_link}}_{a}({v})=av$ (this is the case where $v=z$ in Figure 1). We create a new leaf for $aS$ as a new child of $av$.
In both cases, we add new Weiner links from the nodes in the path between $S$ and $v$ ($S$ inclusive and $v$ exclusive) to the new leaf $aS$ with the character $a$.
Lemma 5 ([17, 2]).
The number of nodes visited during the update from $\mathsf{STree}(S)$ to $\mathsf{STree}(aS)$ is amortized $O(1)$, and the total time cost for updating $\mathsf{STree}(S)$ into $\mathsf{STree}(aS)$ is amortized $O(\log\sigma)$.
The $\log$ factor in the update time is the cost for accessing each Weiner link that is labeled by the new character $a$.
3.3 Computing $\mathsf{NF^{+}}(aS)\triangle\mathsf{NF^{+}}(S)$
Suppose that we have built $\mathsf{STree}(S)$ for string $S$ of length $n$ augmented with soft and hard Weiner links, and that we have computed the list $\mathsf{NF^{+}}(S)$. For each node $u$ of $\mathsf{STree}(S)$ with $\phi_{S}(u)\geq 1$, each element $(u,\phi_{S}(u))\in\mathsf{NF^{+}}(S)$ is stored in $u$. By Lemma 4, we can store $\mathsf{NF^{+}}(S)$ for all nodes $u$ of $\mathsf{STree}(S)$ in total $O(n)$ space.
We update $\mathsf{STree}(S)$ to $\mathsf{STree}(aS)$ with new character $a$ using Weiner’s algorithm, as described above.
By the definition of NF, we obtain the following:
Lemma 6.
For any characters $a,b\in\Sigma$, $(a,b)\in\Phi_{S}(u)$ iff
  • (1)
    $u$ is an internal node of $\mathsf{STree}(S)$,
  • (2)
    $u$ has a soft Weiner link ${\mathsf{W\_link}}_{a}({u})$ that points to a leaf, and
  • (3)
    $u$ has an out-edge that begins with $b$ and leads to a leaf.
Proof.
Suppose $(a,b)\in\Phi_{S}(u)$ holds. Then, by definition, (1) $u$ is a repeat in $S$. Also, $\mathsf{occ}_{S}(au)=\mathsf{occ}_{S}(ub)=\mathsf{occ}_{S}(aub)=1$. By Lemma 1 and by our assumption that $S$ terminates with a unique end-marker \$currency-dollar\\$\$, $u$ is an internal node in $\mathsf{STree}(S)$. Thus, (3) $u$ has an out-edge that begins with $b$ and leads to a leaf. It is clear that $u$ has a Weiner link with $a$. Assume for a contrary that this Weiner link is hard. Then $au$ must be a node in $\mathsf{STree}(S)$. Since $\mathsf{occ}_{S}(au)=1$, $au$ has to be a leaf. But this contradicts that $\mathsf{occ}_{S}(aub)=1$. Thus, (2) ${\mathsf{W\_link}}_{a}({u})$ is a soft Weiner link from $u$. Since $\mathsf{occ}_{S}(au)=1$, this Weiner link points to a leaf.
Suppose (1), (2), and (3) all hold. Then, it is clear that $\mathsf{occ}_{S}(au)=\mathsf{occ}_{S}(ub)=1$ and $\mathsf{occ}_{S}(u)\geq 2$. Assume for a contrary that $\mathsf{occ}_{S}(aub)=0$. Then, since $\mathsf{occ}_{S}(au)=1$, $au$ has to be a leaf. This contradicts that $u$ has a soft Weiner link ${\mathsf{W\_link}}_{a}({u})=aux$ with some non-empty string $x\in\Sigma^{+}$. Hence $\mathsf{occ}_{S}(aub)>0$. Since $\mathsf{occ}_{S}(au)=1$, we have $\mathsf{occ}_{S}(aub)=1$. Therefore, $(a,b)\in\Phi_{S}(u)$. ∎
By the definition of NF, it suffices for us to check the following types of nodes (recall Section 2.3):
  1. (i)
    Existing maximal repeat $u$ in $aS$ that obtains a new left-extension $a$ with $\mathsf{occ}_{aS}(aub)=\mathsf{occ}_{aS}(au)=1$, and has an existing right-extension $b$ with $\mathsf{occ}_{aS}(ub)=1$.
  2. (ii)
    New maximal repeat $w$ in $aS$ that has left-extension $d$ and right-extension $c$ with $\mathsf{occ}_{aS}(dwc)=\mathsf{occ}_{aS}(dw)=\mathsf{occ}_{aS}(wc)=1$.
  3. (iii)-1
    Existing maximal repeat $v$ in $aS$ that has a left-extension $a$ with $\mathsf{occ}_{S}(av)=1$ and $\mathsf{occ}_{aS}(av)>1$.
  4. (iii)-2
    Existing maximal repeat $y$ in $aS$ that has a right-extension $e$ with $\mathsf{occ}_{S}(ye)=1$ and $\mathsf{occ}_{aS}(ye)>1$.
Refer to caption Refer to caption
Figure 2: A snapshot of Weiner’s suffix tree construction algorithm in the case where $ax$ is a leaf. If the leaf $x$ is a child of $v$, then $(a,c)\in\Phi_{S}(v)\setminus\Phi_{aS}(v)$. Let $d$ be a unique character that immediately precedes $ax$ in $S$. Then $(d,c)\in\Phi_{aS}(av)\setminus\Phi_{S}(av)$. If $daz$ in an implicit node on the edge leading to the leaf $dax$, then $(d,e)\in\Phi_{S}(az)\setminus\Phi_{aS}(az)$.
3.4 Dealing with Type (i) node
Here we show that the number of nodes in Type (i) is at most one per iteration. See also Figure 1 for an illustration.
Lemma 7.
If $(u,\phi_{aS}(u))$ is a new element in $\mathsf{NF^{+}}(aS)$ for an existing node $u$, then $u$ is the parent of the leaf representing $S$.
Proof.
Assume for a contrary that $(u^{\prime},\phi_{aS}(u^{\prime}))\in\mathsf{NF^{+}}(aS)\setminus\mathsf{NF^{+}}(S)$ for an existing node $u^{\prime}$ that is not the parent of the leaf for $S$, namely $u^{\prime}\neq u$. By assumption, $\mathsf{occ}_{aS}(au^{\prime})=1$ and $\mathsf{occ}_{S}(au^{\prime})=0$. This means that $u^{\prime}$ is a node on the path $P$ between $v$ and the leaf for $S$, exclusive. Since $u^{\prime}$ is a non-parent ancestor of the leaf for $S$, we have $\mathsf{occ}_{S}(u^{\prime}b^{\prime})\geq 2$, where $b^{\prime}$ is the first character of the out-edge of $u^{\prime}$ on the path $P$. By Lemma 3, $a$ is the unique paring character for $b^{\prime}$, namely $(a,b^{\prime})\in\Phi_{aS}(u^{\prime})$. However, since $\mathsf{occ}_{S}(u^{\prime}b^{\prime})\geq 2$, we have $(a,b^{\prime})\notin\Phi_{aS}(u^{\prime})$, a contradiction. ∎
3.5 Dealing with Type (ii) node
Since Weiner’s algorithm inserts at most one new internal node per iteration, the number of nodes in Type (ii) is at most one per iteration. Let $av$ be the parent of the new leaf representing $aS$. Then, there exists a unique character $d$ that immediately precedes $ax$ in $S$. The following lemma shows the only case where the node $av$ obtains a new pair of characters in $\Phi_{aS}(av)$. See also Figure 2 for an illustration.
Lemma 8.
Let $c$ the first character of the edge from $av$ to $ax$. Then $(d,c)\in\Phi_{aS}(av)$ only if the node $ax$ is a leaf, where $ax$ is the destination node of the Weiner link from $v$ labeled $a$.
Proof.
Since $avc$ is a prefix of $ax$, $davc$ is a prefix of $dax$. Thus, due to Lemma 3, $(d,c)$ is the unique character pair for the character $c$ that can possibly be a member of $\Phi_{aS}(av)$.
If $ax$ is not a leaf, then $\mathsf{occ}_{aS}(avc)\geq\mathsf{occ}_{aS}(ax)>1$, implying $(d,c)\notin\Phi_{aS}(av)$. ∎
3.6 Dealing with Type (iii)-1 node
Here we show that the number of nodes in Type (iii)-1 is also at most one per iteration.
Lemma 9.
Let $v$ be the lowest ancestor of leaf $S$ in $\mathsf{STree}(S)$ such that ${\mathsf{W\_link}}_{a}({v})$ is defined. Then $v$ is the only possible node in Type (iii)-1.
Proof.
Suppose that $v$ is a node in Type (iii)-1, which implies that $\mathsf{occ}_{S}(av)=1$ and $\mathsf{occ}_{aS}(av)>1$. Let $ax$ be the destination node of the soft Weiner link ${\mathsf{W\_link}}_{a}({v})$. Then $av$ is a proper prefix of $ax$, and $av$ is an implicit node in $\mathsf{STree}(S)$ before the update. Since $\mathsf{occ}_{S}(av)=1$, we have $\mathsf{occ}_{S}(ax)=1$ which means that $ax$ is a leaf. Since the right-end of $S$ terminates with \$currency-dollar\\$\$, $x$ is also a leaf.
Consider the case where $x$ is a child of $v$. Then $(a,c)\in\Phi_{S}(v)$, where $c$ is the first character of the out-edge of $v$ leading to $x$ (see also the left diagram of Figure 2). Since $\mathsf{occ}_{aS}(av)>1$, $av$ becomes a branching node in $\mathsf{STree}(aS)$. Thus $(a,c)\notin\Phi_{aS}(v)$ holds for node $v$.
Let $v^{\prime}$ be any node in the path between $v$ and $z$ exclusive. Note that those nodes $v^{\prime}$ are only the other nodes than $v$ such that $\mathsf{occ}_{S}(av^{\prime})=1$ and $\mathsf{occ}_{aS}(v^{\prime})>1$. However, since their descendant $v$ is an internal node, $(a,c^{\prime})\notin\Phi_{aS}(v^{\prime})$ where $c^{\prime}$ is the only possible paring character for $a$ with node $v^{\prime}$.
Thus $v$ is the only possible node in Type (iii)-1. ∎
3.7 Dealing with Type (iii)-2 node
Here we show that the number of nodes in Type (iii)-2 is also at most one per iteration.
Lemma 10.
Let $z$ be the lowest ancestor of leaf $S$ in $\mathsf{STree}(S)$ such that ${\mathsf{W\_link}}_{a}({z})$ is hard. Then $az$ is the only possible node in Type (iii)-2.
Proof.
Suppose that $az$ is a node in Type (iii)-2, which implies that $\mathsf{occ}_{S}(aze)=1$ and $\mathsf{occ}_{aS}(aze)>1$ for a character $e$. Clearly node $az$ has an out-edge that begins with this character $e$. Since $\mathsf{occ}_{S}(aze)=1$, this edge leads to a leaf in $\mathsf{STree}(S)$. Moreover, this leaf is $ax$. Let $d$ be a unique character that immediately precedes $ax$ in $S$. Then, $(d,e)\in\Phi_{S}(az)$ (see also the left diagram of Figure 2).
Consider the case where $av$ becomes a branching node in $\mathsf{STree}(aS)$ after the update. Then, $av$ becomes a child of $az$, meaning that $\mathsf{occ}_{aS}(aze)>1$. Thus $(d,e)\notin\Phi_{aS}(az)$ for node $az$. Since Weiner’s algorithm inserts at most one new branching node, $az$ is the only node in Type (iii)-2. ∎
3.8 Complexities
Theorem 1.
For a string $S$ of length $n$ and a character $a\in\Sigma$, Weiner’s algorithm solves Problem 1 of online NF computation in $O(\log\sigma)$ amortized time per character using $O(n)$ space.
Proof.
It is known that Weiner’s algorithm finds the nodes $u$, $v$, and $z$ in $O(1)$ amortized time for a new character $a$ that is prepended to $S$ (Lemma 5). It follows from our previous discussions that given the nodes $u$, $v$, and $z$, we can compute $\mathsf{NF^{+}}(aS)\triangle\mathsf{NF^{+}}(S)$ in $O(\log\sigma)$ time, where the $\log\sigma$ factor is the cost to access the Weiner links.
By Lemma 2 the number of soft and hard Weiner links is $O(n)$. Since the suffix tree takes $O(n)$ space, the total space requirement is $O(n)$. ∎
Since all positive values of NFs are stored and maintained in the respective nodes of the online suffix tree, we obtain the following:
Corollary 2.
There exists an online data structure that can be maintained in $O(n\log\sigma)$ total time with $O(n)$ space, and can answer Single-NF queries in $O(m\log\sigma)$ time, where $m$ is the query pattern length.
By maintaining a list for $\mathsf{NF^{+}}(S)$ in an online manner, we obtain the following:
Corollary 3.
There exists an online data structure that can be maintained in $O(n\log\sigma)$ total time with $O(n)$ space, and can report all answers to the All-NF problem in output optimal $O(|\mathsf{NF^{+}}(S)|)$ time.
Guo et al. [8] also considered a version of the All-NF problem, named All-NF-Extract, that requires us to compute the product $|u|\cdot\phi_{S}(u)$ for all pairs $(u,\phi_{S}(u))\in\mathsf{NF^{+}}(S)$. Guo et al. [8] proposed an offline algorithm that works in $O(n\log\delta)$ time for All-NF-Extract for integer alphabets, where $\delta$ denotes the substring complexity [10] of the input string. Our online algorithm can readily be extended to the All-NF-Extract problem:
Corollary 4.
There exists an online data structure that can be maintained in $O(n\log\sigma)$ total time with $O(n)$ space, and can report all answers to the All-NF-Extract problem in output optimal $O(|\mathsf{NF^{+}}(S)|)$ time.
Acknowledgements
This work is supported by JSPS KAKENHI Grant Numbers JP20H05964, JP23K24808, JP23K18466.
The author thanks Takuya Mieno for fruitful discussions.